User Tools

Site Tools


operations:handover

Handover notes

Please use this page to write up any notes on anything that needs to be passed on to the next observer or should go into the end of experiment message.

Notes should be placed below the horizontal line as a new blog entry.

Continuing problems should be copied to the Current Issues page.

**Template for Handover notes:**

Be sure to add any specific notes about the experiment if required, for example “Special back end configuration for x reason”. Put as much detail as you can - the more the better.

It is ESSENTIAL to put your name after EVERY entry. Copy and paste the template into the new handover notes

====== Experiment Name ======

**Hobart 12m**

Disk VSN: abc-123 | Data volume at beginning: 2.34 GB

   * 1300UT a comment from me (user)

**Katherine 12m**

Disk VSN: | Data volume at beginning:  GB

   * 

**Yarragadee 12m**

Disk VSN: | Data volume at beginning:  GB

   * 1300UT Experiment started okay (JMc).

Format for entering a comment It is important to stick to this formula - there is a good reason we need this information, so please include it!

Time stamp in UT * Comment * (Your Name)

For Example:

20:00UT: Wind stows, Missed scans 012-3456 to 012-3457 (Jim)

20:50UT: Mark5 lost connection, had to restart Mark5 and reconfigure DBBC, called Jamie on call to fix, missed scans 012-3456 to 012-4567 (Jim)

Again - as much detail as you can. It is important to remember that these notes will be sent to the correlator, so please make sure they make sense to others.


RD2001

Hobart 26m

Disk VSN: UVLBI+83/6000/1024 | Data volume at beginning: 1.154 GB

  • 1800UT - Experiment started ok (Prad).
  • 2015UT - The system switched off. Had to turn on the UPS and CPU. Experiment wasn't affected (Prad).
  • 2134UT - FS crashed. Restarted FS. Scan 015-2134 potentially affected (Prad).
  • Power failure at MtP. Dish came back on source at 2217UT.
  • 70K stage temperature values shoot up from time to time. Display/sensor issue?
  • Power failure at around 2230UT. Restarted system at 2243UT. Scans 015-2229 and 015-2242 affected.
  • 2250UT - Wind stowed is what the live page indicated. Back on at 2259UT. Scan 015-2250 missed.
  • 0733UT mk5cn: time-out, connection closed occasionally happened after a few “recording stopped”. Recording started normally for the subsequent scans.
  • 1606UT - looks like it fails to connect to mk5, auto correlation plot shows correct timestamp but scan number “-alpha-”, warning “mk5cn: error opening, mark 5 probably not running, see above for error” without error. disk_pos doesn't update. One of the Mk5hb light flickered, ip switching mk5hb made one of the light on module flickers. (Lim)
  • Problem not solved till the end of schedule. Scans very likely affected. (Lim)
2020/01/15 17:33 · Pradyumna Kiran Kummamuru
OHG123

Hobart 26m

Disk VSN: HOB+0015 | Data volume at beginning: 2217.943 GB

  • Eremotectl not working for Ho. Using stream_log on ops8 (Patrick)
  • 1730UT - Experiment started okay (Patrick)
  • 1030UT to 1036UT - ALARM: error m5 -104 mk5cn: time-out, connection closed. It seems it was an on and off internet connection then it fixed. Affected scan: 015-1033 (Simin).

Yarragadee 12m

Disk VSN: TR-00011 | Data volume at beginning: 0.256 GB

  • 1730UT - Experiment started okay (Patrick)
2020/01/14 17:10 · Pradyumna Kiran Kummamuru
R1929

Hobart 26m

Disk VSN: HOB+0015 | Data volume at beginning: 0 GB

  1. 1700 UT Experiment started OK (Prad)
  2. 0606 UT eRemote not working on any computers, so just using newsmerd, etc. (Ross)
2020/01/13 19:44 · Pradyumna Kiran Kummamuru
R4928

Yarragadee 12m

Disk VSN: MED-0027 | Data volume at beginning: 0.513 GB

  1. 1832 UT Experiment started OK (mas)
2020/01/09 18:48 · Patrick Yates
R1928

Yaragadee 12m

  16:50 Experiment Started OK (Earl).
2020/01/07 16:53 · Earl Lester · 0 Comments
R4927

Yarragadee 12m

Disk VSN: USN+0042 | Data volume at beginning: 0.00 GB

  • Pointing check kept failing, the wind speed was quite high though.
  • 1830UT Experiment started okay (Lucas).
2020/01/02 21:37 · Tiege McCarthy · 0 Comments
/home/www/auscope/opswiki/data/pages/operations/handover.txt · Last modified: 2018/10/29 16:27 by Jesse Swan